OpenStudy (anonymous):

13k + 5 = k - 7 All Work

OpenStudy (diyadiya):

13k-k=-7-5 12k=-12 k=-1

OpenStudy (saifoo.khan):

13k - k = -7 -5 12k = -12 k = -12/12 k = -1